OpenStudy (anonymous):

solve by factoring 5x^3=45

OpenStudy (anonymous):

5x^(3)=45 Divide each term in the equation by 5. (5x^(3))/(5)=(45)/(5) Simplify the left-hand side of the equation by canceling the common terms. x^(3)=(45)/(5) Simplify the right-hand side of the equation by simplifying each term. x^(3)=9 Take the cube root of both sides of the equation to eliminate the exponent on the left-hand side. x=~3:(9)_x<Z>APPR<z>2.08

OpenStudy (anonymous):

oh ok,well the options i have are 3,-3,0,10

OpenStudy (anonymous):

3 and -3 would only work if it was \[x^{2}\] not \[x^{3}\]

OpenStudy (anonymous):

have a look

OpenStudy (anonymous):

3

OpenStudy (anonymous):

did u received the atrtachment

OpenStudy (anonymous):

reply me